In a crank and slotted lever mechanism, the distance between the fixed center is 150mm and the driving crank is 75mm long. Determined the ratio of the time taken on the cutting and return strokes .


1
Expert's answer
2021-02-15T05:41:52-0500

Solution:-

In a crank and slotted lever quick return mechanism, the distance between the fixed centers is 150 mm and the driving crank is 75 mm long. Determine the ratio of the time taken on the cutting and return strokes:-

2.0

Explanation:-

The schematic of quick return mechanics distance between fixed centers-

AB = 150 mm



Length of driving crank,

BC = 75 mm

 In triangle ABC,

"cos\theta=\frac{BC}{AB}=\frac{75}{150}=0.5 \implies \theta=60^0"

Since cutting of metal happens to be in forward stroke and angle rotated by crank BC during cutting

"\alpha=360-2\theta=360-120=240"

 Ratio of cutting stroke time to return stroke time,

"=\frac{\alpha}{\beta}=\frac{240}{120}=2"